OpenStudy (loser66):

What is wrong? Solve \(13x \equiv 6 (mod 15) Please, help

OpenStudy (loser66):

That is 13x -15y = 6 gcd (13, 15) =1 and 1 | 6, hence there is a solution 15 = 13 +2 , then 2 = 15 -13 13 = (7) 2 + 1, then 1 = 13 -(7) 2 = 13-(7) (15-13) = (8) 13 -(7) 15 multiple both sides by 6 13(48) - 15(42) =6 Hence x = 48 mod 15, or \(x \equiv 3 (mod 15)\)

OpenStudy (loser66):

Now, crazy thing comes. if I take x = 18, then \(x\equiv 3 (mod 15)\) then x =18 is one of the solution, right? Plug back to original one 13*18 = 234 \(\equiv 9 (mod 15)\) not 6 as required. :(

OpenStudy (loser66):

@superdavesuper

OpenStudy (loser66):

@IrishBoy123

OpenStudy (loser66):

But if I stop at 2 = 15 -13, I am ok, since 6 = 15 (3) + 13( -3) Hence \(x \equiv -3\equiv 12 (mod 15)\) Test back , take x =12, then 13 * 12 = 156 \(\equiv 6 (mod 15)\) I don't know what is wrong with the first one.

OpenStudy (superdavesuper):

how about simple typo... " 15 = 13 +2 , then 2 = 15 -13 13 = (7) 2 + 1, then..." second line should be 13 = (6)2 + 1 going down dat path, u will get 12 (mod 15) as the right ans :D

OpenStudy (loser66):

damn!! silly mistake. Thank you so much